OpenStudy (anonymous):

describe the relationship between distance and acceleration.how would an increase or decrease in one value affect the other

OpenStudy (aayushi.somani):

\[a=\frac{ d ^{2} x}{ d ^{2} t}\] where x= distance, a= acceleration and t= time

OpenStudy (anonymous):

x(t) = a int dx(t) = v(t) = ax(t) + b int v(t) = a(t) = (ax(t)^2)/2 +bx +c